# Surrogate model for Bayesian optimal experimental design in chromatography

### v1

@@ 4.1 Problem formulation
In our numerical simulations, we model a chromatography process for two components, i.e., $N_{c}=2$. Moreover, we assume that our measurements are obtained on an equidistant temporal grid with $N_{s}$ nodes and the measurement data is contaminated with a normally distributed noise with standard deviation $\sigma$.
We utilize a conservative prior model composed of independent uniform distributions for every component of the parameter $\theta$. That is, for every parameter $\theta_{i}$, we define a uniform prior in each interval $\left[\alpha_{i},\beta_{i}\right]\subset\mathbb{R}_{+}$, i.e., $\theta_{i}\sim\mathcal{U}\left(\left[\alpha_{i},\beta_{i}\right]\right)$. Furthermore, we can restrict the domain to the Cartesian product of the intervals $\Theta=\bigtimes_{i=1}^{4}\left[\alpha_{i},\beta_{i}\right]$. We assume that these intervals can be reliably specified from independent information, e.g. a rough preliminary experiment. We note that similar methodology was utilized in [61].
@@ 4.2 Simulations
In the spirit of [11], we specified the true data-generating parameters as $b_{1}=0.05$, $b_{2}=0.10$, $Q_{s}=10$. Moreover, we set $N_{tp}=70$. The standard deviation of noise was $\sigma=0.05$, and other parameters satisfied $\Upsilon=10$ and the constant $F=1.5$. The intervals of the prior distribution and for the design space are shown in the tables 1 and 2, respectively.
$Parameter$ Lower Bound Upper Bound Real parameter
$b_{1}$ 0.02 0.08 0.05
$b_{2}$ 0.03 0.17 0.1
$Q_{s}$ 8 11 10
$N_{tp}$ 50 180 70
@@
$Parameter$ Lower Bound Upper Bound
$\tau^{inj}$ 0.05 3
$c^{Feed}$ 1 15
@@
In all cases we estimated the utility function with a high-resolution forward mapping and with PSLI for $N=1105$ training nodes. Every training process was carried out on an equidistant grid over the design space with 14 nodes in each direction. The average times needed for train the surrogate model is given in Table 3. We compare the times invested in training the surrogate model by using the high resolution model. In the case of the utility function, we can see how the training time of the surrogate is one order of magnitude smaller than the brute-force evaluation with real model. The time needed to calculate a parameter chain of length 80 000 with MCMC is three orders of magnitude smaller than that using the real model. The utility functions are presented in the Figure 5.
@@
To validate the utility function results, we selected 6 design nodes (L, M, N, R, T, and X) and computed MCMC samples using both the true model and the surrogate model at these points. These nodes were chosen to facilitate comparisons across rows and columns, demonstrating that the optimal posterior distribution does not always directly correlate with $\tau^{inj}$ and $c^{Feed}$. The posterior sampling was carried out with the DRAM algorithm [21], comprising 80 000 simulations with the first 30 000 samples discarded to mitigate the burn-in effect. The results are illustrated in figures 6, 7, and 8.
@@
Intuition for the observations above can be drawn from the figures 2, 3, and 4, where the synthetic data was generated. For $N_{s}=8$ and point $L$ we can see how in this narrow peak, only the second temporal node contributes for the component $c_{1}$. It implies that even that the measurement has larger magnitude than node $N$ or $R$, then the node $L$ could not be able to capture the information with perturbations in the parameter space. The same phenomena explains why $X$ and $M$ are the best and the worst design variables respectively.
@@
Figure 5: Utility functions (time injection $\tau^{inj}$ in $x$ axis and concentration $c^{Feed}$ in $y$ axis). Left: surrogate models for 8, 15, and 20 temporal nodes of the sensor respectively. Middle: the utility function with the true model. Right: absolute errors.
Figure 6: Samples of MCMC. In the left surrogate models for $N_{s}=8$. In the right the samples with the true model
Figure 7: Samples of MCMC. In the left surrogate models for $N_{s}=15$. In the right the samples with the true model
Figure 8: Samples of MCMC. In the left surrogate models for $N_{s}=20$. In the right the samples with the true model
@@
Table 3: Average time of evaluations in seconds. Row one: the forward mapping $G(\theta,d)$, row two: the utility function $U(d)$ at one design point, row three: MCMC parameter chain of length 80 000.
@@ 5 Conclusions
The surrogate model, based on PSLI, significantly reduced evaluation time compared to solving the PDE system with the Koren scheme. Moreover, it reproduced evaluations from the original model to a good approximation, as confirmed through simulations of the expected utility functions with different designs. This observations was enforced by considering accuracy in temporal profile prediction and simulation of posterior ensembles to compare concentrations. We note that other methods tested, such as polynomial chaos expansion, failed to produce satisfactory results due to high gradient zones in the concentration profiles.

### v2

@@ 4.1 Problem formulation
In our numerical simulations, we assume that our measurements are obtained on an equidistant temporal grid with $N_{s}$ nodes and the measurement data is contaminated with a normally distributed noise with standard deviation $\sigma$.
We utilize a conservative prior model composed of independent uniform distributions for every component of the parameter $\theta$. That is, for every parameter $\theta_{i}$, we define a uniform prior in each interval $\left[\alpha_{i},\beta_{i}\right]\subset\mathbb{R}_{+}$, i.e., $\theta_{i}\sim\mathcal{U}\left(\left[\alpha_{i},\beta_{i}\right]\right)$. Furthermore, we can restrict the domain to the Cartesian product of the intervals $\Theta=\times_{i=1}^{4}\left[\alpha_{i},\beta_{i}\right]$. We assume that these intervals can be reliably specified from independent information, e.g. a rough preliminary experiment. We note that a similar methodology was utilized in [63].
@@ 4.2 Simulations
In the spirit of [11], we specified the true data-generating parameters as $b_{1}=0.05$ [L/mol] , $b_{2}=0.10$ [L/mol], $Q_{s}=10$ [mol/L]. Moreover, we set $N_{tp}=70$. The standard deviation of noise was $\sigma=0.05$ [mol/L], and other parameters satisfied $\Upsilon=10$ and the constant $F=1.5$. The intervals of the prior distribution and for the design space are shown in the tables 1 and 2, respectively.
$Parameter$ Lower Bound Upper Bound Real parameter Units
$b_{1}$ 0.02 0.08 0.05 L/mol
$b_{2}$ 0.03 0.17 0.1 L/mol
$Q_{s}$ 8 11 10 mol/L
$N_{tp}$ 50 180 70 -
@@
$Parameter$ Lower Bound Upper Bound Units
$\tau^{inj}$ 0.05 3 -
$c^{Feed}$ 1 15 mol/L
@@
In all cases we estimated the utility function with a high-resolution forward mapping and with PSLI for $N=1105$ training nodes. Every training process was carried out on an equidistant grid over the design space with 14 nodes in each direction. The average times needed to train the surrogate model are given in Table 3. We compare the times invested in training the surrogate model by using the high resolution model. In the case of the utility function. We can see how the training time of the surrogate is one order of magnitude smaller than the brute-force evaluation of the utility function with the real model. The time needed to calculate a parameter chain of length 80 000 with MCMC is three orders of magnitude smaller than that using the real model. The utility functions are presented in Figure 2.
@@
To validate the utility function results, we selected 6 design nodes (L, M, N, R, T, and X) and computed MCMC samples using both the true model and the surrogate model at these points. These nodes were chosen to facilitate comparisons across rows and columns, demonstrating that the optimal posterior distribution does not always directly correlate with $\tau^{inj}$ and $c^{Feed}$. The posterior sampling was carried out with the DRAM algorithm [21], comprising 80 000 simulations with the first 30 000 samples discarded to mitigate the burn-in effect. The results are illustrated in Figures 6-11.
@@
Intuition for the observations above can be drawn from the Figures 3, 4, and 5, where the synthetic data was generated. For $N_{s}=8$ and point $L$ we can see how in this narrow peak, only the second temporal node contributes for the component $c_{1}$. It implies that even though the measurement has larger magnitude than node $N$ or $R$, the node $L$ could not be able to capture the information with perturbations in the parameter space. The same phenomena explains why $X$ and $M$ are the best and the worst design variables respectively.
@@
Figure 2: Utility functions. Left: true models for 8, 15, and 20 temporal nodes of the sensor respectively. Middle: the utility function with the surrogate model. Right: absolute errors.
@@
Figure 6: Samples of MCMC, $\theta_{1}$ vs $\theta_{2}$. In the left surrogate models for $N_{s}=8$. In the right the samples with the true model.
Figure 7: Samples of MCMC, $\theta_{3}$ vs $\theta_{4}$. In the left surrogate models for $N_{s}=8$. In the right the samples with the true model.
Figure 8: Samples of MCMC, $\theta_{1}$ vs $\theta_{2}$. In the left surrogate models for $N_{s}=15$. In the right the samples with the true model.
Figure 9: Samples of MCMC, $\theta_{3}$ vs $\theta_{4}$. In the left surrogate models for $N_{s}=15$. In the right the samples with the true model.
Figure 10: Samples of MCMC, $\theta_{1}$ vs $\theta_{2}$. In the left surrogate models for $N_{s}=20$. In the right the samples with the true model.
Figure 11: Samples of MCMC, $\theta_{3}$ vs $\theta_{4}$. In the left surrogate models for $N_{s}=20$. In the right the samples with the true model.
@@
Table 3: Average time of evaluations in seconds. Row one: the forward mapping $G(\theta,d)$, row two: the utility function $U(d)$ with 10000 evaluations of $G(\theta,d)$ at one design point, row three: MCMC parameter chain of length 80 000.
@@ 5 Conclusions
The surrogate model, based on PSLI, significantly reduced evaluation time compared to solving the PDE system with the Koren scheme. Moreover, it reproduced evaluations from the original model with a small approximation error, as confirmed through simulations of the expected utility functions with different designs. This observation was enforced by considering accuracy in temporal profile prediction and simulation of posterior ensembles to compare concentrations. We note that other methods tested, such as polynomial chaos expansion, failed to produce satisfactory results due to high gradient zones in the concentration profiles.